# REDS: Resource-Efficient Deep Subnetworks for Dynamic Resource Constraints

### v2

@@ 3.4 REDS fine-tuning
100% 93.83±$\pm 0.22$ 93.38±$\pm 0.45$ 93.34±$\pm 0.21$ 93.19 ±$\pm 0.18$ 94.87 ±$\pm 0.33$ 94.46±$\pm 0.08$ 94.35 ±$\pm 0.21$ 94.05 ±$\pm 0.14$
75% 93.37±$\pm 0.34$ 93.18±$\pm 0.21$ 93.03±$\pm 0.26$ 92.85 ±$\pm 0.47$ 94.27 ±$\pm 0.08$ 94.32±$\pm 0.13$ 94.18 ±$\pm 0.17$ 93.76 ±$\pm 0.02$
50% 93.41±$\pm 0.67$ 92.12±$\pm 0.24$ 92.26±$\pm 0.50$ 91.50 ±$\pm 0.10$ 94.11 ±$\pm 0.26$ 94.17±$\pm 0.01$ 94.00 ±$\pm 0.25$ 93.08 ±$\pm 0.06$
25% 91.46±$\pm 0.80$ 89.64±$\pm 0.76$ 88.59±$\pm 1.69$ 85.14 ±$\pm 1.07$ 93.80 ±$\pm 0.14$ 93.20±$\pm 0.19$ 93.17 ±$\pm 0.73$ 92.11 ±$\pm 0.42$
Table 1: Test set accuracy [%] of training S and L depth-wise separable convolutional architectures (DS-CNN) from (Zhang et al., 2017): training a network of each size from scratch ("Scratch"), conversion from a pre-trained network using two knapsack versions ("Knapsack BU" and "Knapsack TD"), and training REDS structure from scratch ("REDS training"). Reported results from three independent runs. The accuracy of each 100 % network reported by Zhang et al. (2017) is listed in the header row.
4 REDS Performance
4.1 Evaluation setup
We evaluate the performance of REDS, which comprises four nested subnetworks obtained by multiplying the full model MACs (100%) by the predefined constraints percentages (25%, 50%, and 75%), on three datasets: Google Speech Commands (Warden, 2018), FMNIST (Xiao et al., 2017), CIFAR10 (Krizhevsky et al., 2009). In addition, two nested subnetworks are evaluated on the Visual Wake Words dataset (Chowdhery et al., 2019) for 50% of MACs and 200 KB of peak memory usage as constraints. We measure the subnetworks’ peak memory usage using the open-source tf-lite tools (Liberis et al., 2021), assuming the temporary buffer space is utilized for storing both input and output buffers (Chowdhery et al., 2019). REDS is evaluated on three pre-trained Google Speech Commands architectures (DNN, CNN, and DS-CNN) of two sizes (S and L) each introduced in Zhang et al. (2017), and on an ImageNet (Deng et al., 2009) pretrained MobileNetV1 (Howard et al., 2017) architecture fine-tuned on the Visual Wake Words dataset. For CIFAR10 and FMNIST REDS is tested on two pre-trained DS-CNN size S architectures. DNN is a 2-layer fully-connected architecture with layer width 144 (S) and 436 (L). CNN is an architecture with two convolutional layers followed by 3 fully-connected layers. The widths of convolutional layers are 28 and 30 for the S architecture and 60 and 76 for the L architecture, respectively. DS-CNN and MobileNetV1 are composed of a standard convolutional layer followed by several blocks of depth-wise and point-wise convolutional layers (see Fig. 2). The former has 4 blocks for the network of size S, and 5 blocks for the network of size L with a layer width of 64 and 276 respectively. The latter includes 13 blocks with an initial layer width of 32 that progressively increases to 1024 with the model’s depth. Each convolutional layer in the CNN, DS-CNN and MobileNetV1 architectures is followed by a batch normalization layer (Ioffe and Szegedy, 2015).
4.2 REDS empirical evaluation
The main paper presents the results for DS-CNN architectures (S and L) on Google Speech Commands and MobileNetV1 (0.25x) on Visual Wake Words. The fine-tuning hyperparameters and additional results for other architectures and datasets are available in the extended Appendix C. The REDS structure for Google Speech Commands used in the main paper comprises only four nested subnetworks, however a larger number of subnetworks does not degrade the accuracy of REDS fine-tuned on Google Speech Commands (see Fig. 10). REDS was tested on tasks of varying difficulty (FMNIST and CIFAR10) without accuracy degradation compared to training from scratch. In contrast to the state-of-the-art optimizers for edge devices, such as $\mu$NAS (Liberis et al., 2021), REDS demonstrate a faster solution search time, taking only a few minutes for DS-CNN and hours for MobileNetV1 as opposed to days by $\mu$NAS.
Figure 4: DS-CNN S and L few-shots finetuning on Google Speech Commands (Warden, 2018) with MACs percentages as constraints. The subnetworks obtained from the BU Knapsack formulation exhibit faster recovery in accuracy compared to the TD Knapsack ones.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: DS-CNN S and L few-shots finetuning on Google Speech Commands (Warden, 2018) with MACs percentages as constraints. The subnetworks obtained from the BU Knapsack formulation exhibit faster recovery in accuracy compared to the TD Knapsack ones.
@@ 4.2 REDS empirical evaluation
We observe that the accuracies of each submodel for the same percentage of MACs are similar, even for small S architectures with only 25 % of MACs. The accuracies of the full models, highlighted in bold in the top row are computed using pre-trained models from Zhang et al. (2017) and closely match our experimental results. The performance difference between the BU and TD heuristics is largely insignificant due to the fine-tuning step with extensive training data. Even though the models are overparameterized and in all cases achieve high performance, low-capacity subnetworks S sliced by the bottom-up approach yield minor performance gains. However in the low-data regime, where only few samples per class are available for fine-tuning, the performance differences between BU and TD heuristics are remarkable. Fig. 5 shows the performance when few-show learning is applied to fine-tune REDS structures. Our empirical findings confirm a consistently superior performance of the BU heuristic compared to the TD alternative. The differences diminish as more samples are used for fine-tuning (Entezari et al., 2023) and there is barely any difference if full-finetuning is applied using all available training data.
For the MobileNetV1 architecture, the REDS knapsack BU heuristic surpasses the fully-trained uniform pruning in terms of peak memory usage, while maintaining a negligible drop in accuracy (see Table 2). This validates the efficacy of the knapsack BU approach in constraining the peak memory usage within the subnetworks’ architecture. Furthermore, imposing a peak memory usage constraint is essential to facilitate the deployment of the REDS structure within the confines of the available device RAM (see Fig. 7).
@@
Table 2: Comparative analysis of REDS and MobileNet v1 (0.25x) from Chowdhery et al. (2019) on Visual Wake Words. REDS exhibits lower peak memory usage compared to the uniformly pruned fully trained network, with a negligible drop in accuracy.
@@ 5 REDS Optimization for Caches
Embedded ML frameworks like Tensorflow Lite typically store model weight matrices in row-major order. This means, that each row of a weight matrix is stored contiguously in memory. Without the use of sub-networks these weights are also accessed in a contiguous fashion. However, when using subnetworks for model inference, some neurons and their corresponding weights are omitted, resulting in non-contiguous memory access. This effect is illustrated in Figure 1. We now show how by simply adapting the computational graph at compile time, we are able to optimize the computation of REDS subnetworks for devices with a cache memory architecture.
@@ 5.1 Row-major and column-major stores
The calculation of a fully-connected neural network layer during the forward pass is a matrix multiplication $H=\sigma(X_{[m\times b]}^{T}\cdot W_{[m\times n]})$, where $x$ is an input matrix of shape $m\times b$ (input samples $\times$ batch size), $W$ the layer weights matrix of shape $m\times n$ (input samples $\times$ number of neurons) and $\sigma$ the activation function. For simplicity we omit the typically used bias term. During matrix multiplication, like it is for instance implemented in Tensorflow Lite, an inner loop multiplies and sums each element of row $x_{i}$ of $X^{T}$ with column $w_{j}$ of $W$. This column-wise access of row-major ordered weights $w\in W$ may lead to consecutive reads from memory addresses, which are not located close-by depending on the size of the matrix and the use of the REDS subnetworks. However, this can easily be circumvented by changing the matrix multiplication involving a transposed $W$, i.e., $\sigma(X^{T}\cdot W)=\sigma((W^{T}\cdot X)^{T})$, which ensures contiguous memory accesses of the weights. Note that the additional transposition of the resulting matrix introduces no additional computations if the input matrix is a single sample, i.e., $X$ has size $[m\times 1]$. Both $X_{[m\times 1]}$ and $X_{[m\times 1]}^{T}$ are stored equally in memory and, thus, the transposition is redundant in this case. Let us consider the simple example of $X_{[2\times 2]}$ and $W_{[2\times 3]}$ both stored in row-major mode. In the basic matrix multiplication cases $X^{T}\cdot W$, the order of the relative memory access locations of weights $w$ is $0\rightarrow 3\rightarrow 1\rightarrow 4\rightarrow 2\rightarrow 5\rightarrow 0\rightarrow 3\rightarrow\cdot\cdot\cdot$. In the optimized cases $(W^{T}\cdot X)^{T}$ where $W^{T}$ is also stored in column-major form, the order of weights accesses is $0\rightarrow 1\rightarrow 0\rightarrow 1\rightarrow 0\rightarrow 1\rightarrow 2\rightarrow 3\rightarrow\cdot\cdot\cdot$. Depending on the memory and cache setup of a device, this highlights the potential for exploiting cache systems by simply changing the matrix multiplications to the optimized transposed form.
@@ 5.2 Optimizing REDS computational graph
The optimization proposed above can be implemented by assuring that each matrix multiplication within a computational graph of each submodel follows the for cache optimized access pattern and the corresponding weight matrices are stored in column-major order in memory. We now show the effect of using this optimized computational graph by benchmarking different matrix multiplications. To this end, we use a Raspberry Pi Pico board, which features a RP2040 chip based on a dual-core Arm Cortex-M0+ processor architecture with 264kB RAM and 16MB off-chip flash memory (RP2040, 2023). The chip also features execute-in-place (XIP) support, such that the flash memory can be treated like internal memory. These accesses to flash are cached with a 16kB, two way set-associative cache with 8 byte wide cache lines.
Fig. 6 shows the cache effect on the duration of matrix multiplications when we use the optimized computation graph compared to the basic one, i.e., the default way where W is stored in row-major mode and we calculate $X^{T}\cdot W$. We benchmark the matrix multiplication at different splits and use 25%, 50%, 75% and 100% of the neurons in $W$. In Fig. 6 (left) we show the duration of a matrix multiplication with differently shaped weights $W$ times an input matrix $X$ of shape $256\times 4$, where the weights and inputs are 32 Bit wide floating points. Similarly, Fig. 6 (right) shows the duration of a weights matrix multiplication of size $512\times 256$ with the same $X$, but now the weights and inputs are unsigned integers with different widths, i.e., 8, 16 and 32 Bit. We observe in all cases a clear speed-up in matrix multiplication when comparing the optimized and the basic computation approach, which is also evident from Table 3. For all test cases the cache-hit rate is above 97% when using optimized matrix multiplications. Additionally, we see a notable difference between floating point and integer based matrices. This difference is due to the fact that the RP2040 does not feature any floating point arithmetic support and, thus, the cache effect is diminished due to longer duration of basic multiplications. Finally, there is no notable difference in speed-up and cache-hit rate between the different splits.
@@
Figure 7: Peak memory usage analysis of MobileNetV1 processed by the REDS knapsack BU heuristic. Placing constraints on MACs (left), and on both MACs and peak memory usage (right). The peak memory usage constraint enables REDS subnetworks optimized for edge devices.
@@ 6 REDS on Mobile and IoT
We evaluate REDS models on two mobile phones, specifically Xiaomi Redmi Note 9 Pro and Google Pixel 6, and on two IoT devices, namely Arduino Nano 33 BLE Sense and Infineon CY8CKIT-062S2. For the mobile phone evaluation, we employed the official Google TFLite benchmarking tool, which measures the model’s inference time on Android-based mobile devices. On IoT devices we deployed and evaluated the models using the cloud-based platform Edge Impulse (Hymel et al., 2022). We report the number of parameters, accuracy, and latency on the mobiles and on the IoT devices for the full architectures (100% MACs) and the subnetworks architecture (75%, 50% and 25% MACs) found by the BU knapsack solvers. The obtained results are reported in Fig. 8 for the DNN, CNN and DS-CNN architectures S (top row) and L (bottom row) on Google Speech Commands. All results are averages over three runs. TFLite and TFLMicro lack support for runtime adaptation of model weight tensors. We extended the TFLMicro framework to support REDS out of the box [^2] https://github.com/FraCorti/Deep_Subnetworks_for_Dynamic_Resource_Constraints/blob/main/README.md#tensorflow-lite-for-microcontrollers-analysis and measure on-device inference and submodel switching times.
@@
The first left-most column of plots in Fig. 8 shows the relationship between the subnetworks’ MACs and the number of model parameters. All curves are close to linear, yet the parameters of these linear relationships are architecture-specific. The DS-CNN models have the lowest number of parameters, thanks to their more sophisticated architecture, which is more efficient on our dataset compared to the standard convolutional and dense networks. DS-CNN models also yield better accuracy, even for only 25% of MACs, while DNN models perform worst, as can be observed in the second column of plots. This result can be attributed to the successful generalization of the iterative knapsack problem to support depth-wise convolutions (cf.f Appendix A.1). We also compare REDS to several baselines. Recent works Tan and Le (2019) consider model optimization by pruning an equal share of computational blocks in each layer. We use the block selection strategy based on the weights magnitude (i.e., L1) used in Cai et al. (2020), and a random selection. REDS knapsack solution outperforms both baselines achieving higher accuracy for the same percentage of MACs. In addition, the red star in the first two figures shows the performance of the constrained neural architecture search for microcontrollers $\mu$NAS as reported in Liberis et al. (2021). The network architecture uses depth convolutions, yet is more complex than our DS-CNN. $\mu$NAS shows 3% performance improvement over DS-CNN at the cost of 32% higher model size. However, $\mu$NAS took 39 GPU-days to compute the solution, in contrast to minutes for DS-CNN and hours for MobileNetV1 taken by our knapsack solver.
@@
The last two plots on the right show REDS performance on mobile and IoT devices. We observe a difference of three orders of magnitude in the inference times between the two categories. DNN models perform best thanks to the more optimized algorithm and libraries for matrix-matrix multiplication (Goto and Geijn, 2008). All models show a linear relationship between the percentage of MACs and inference time. This empirical evidence validates MACs as a robust predictor of model latency, extending its applicability to mobile devices. We also evaluate REDS against early exit linear classifiers and magnitude-based uniform pruning on the DS-CNN S and L architectures with respect to MACs and test set accuracy, as shown in Fig. 10. The knapsack solution of REDS demonstrates superior performance over both methods, most notably within the domain of lower parameter configurations, specifically the S size.
@@
Figure 9: DS-CNN S (left) and L (right) BU comparison to early exit linear classifiers and subnetworks uniform magnitude (L1) structural pruning.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 9: DS-CNN S (left) and L (right) BU comparison to early exit linear classifiers and subnetworks uniform magnitude (L1) structural pruning.
Figure 10: DS-CNN S (left) and L (right) BU REDS ten subnetwork configuration. A larger number of subnetworks does not degrade the accuracy of REDS.
7 Conclusion, Discussion, Outlook
This paper presents Resource-Efficient Deep Subnetworks (REDS), a novel approach to adapt deep neural networks to dynamic resource constraints typical for mobile and IoT devices. REDS formulates the subnetworks architectures search as an iterative knapsack problem, taking into account the dependencies between layers, MACs and peak memory usage. Moreover, REDS employs neuron permutation invariance to facilitate adaptation to variable resource availability without compromising the models’ accuracy or adding runtime overhead. Notably, REDS ensure the subnetworks’ weights are stored in contiguous memory, enhancing cache optimization and compatibility with hardware-specific enhancements like vector instructions. Experimental evaluation on seven benchmark architectures demonstrates the effectiveness of REDS on mobile and IoT devices and superior performance of current pruning and neural architecture search state of the art methods. We perform a theoretical analysis of the knapsack solution space and prove the worst-case performance bounds for the two heuristic algorithms.
Discussion and Outlook. Our study has some limitations. First, we do not test the efficiency of REDS when combined with quantized models, which is left for future work. Secondly, the support of REDS for specific layers like skip connections is not explored and should be addressed in the future to support further advanced architectures like MobileNet v2. We plan to augment REDS with a task scheduler and deploy REDS in the context of a specific application with dynamically changing resource constraints (Fang et al., 2018).
Acknowledgements
The authors are grateful to Markus Gallacher for his support with the energy efficiency analysis of REDS. Christopher Hinterer and Julian Rudolf contributed to extending the TFL Micro framework to support REDS on Arduino Nano 33 BLE Sense. This research was funded in part by the Austrian Science Fund (FWF) within the DENISE doctoral school (grant number DFH 5). The results presented in this paper were computed using computational resources of HLR resources of the Zentralen Informatikdienstes of Graz University of Technology.

### v3

@@ 4.2 REDS empirical evaluation
100% 93.83±0.22$\pm$ 0.22± 0.22 93.38±0.45$\pm$ 0.45± 0.45 93.34±0.21$\pm$ 0.21± 0.21 93.19 ±0.18$\pm$ 0.18± 0.18 94.87 ±0.33$\pm$ 0.33± 0.33 94.46±0.08$\pm$ 0.08± 0.08 94.35 ±0.21$\pm$ 0.21± 0.21 94.05 ±0.14$\pm$ 0.14± 0.14
75% 93.37±0.34$\pm$ 0.34± 0.34 93.18±0.21$\pm$ 0.21± 0.21 93.03±0.26$\pm$ 0.26± 0.26 92.85 ±0.47$\pm$ 0.47± 0.47 94.27 ±0.08$\pm$ 0.08± 0.08 94.32±0.13$\pm$ 0.13± 0.13 94.18 ±0.17$\pm$ 0.17± 0.17 93.76 ±0.02$\pm$ 0.02± 0.02
50% 93.41±0.67$\pm$ 0.67± 0.67 92.12±0.24$\pm$ 0.24± 0.24 92.26±0.50$\pm$ 0.50± 0.50 91.50 ±0.10$\pm$ 0.10± 0.10 94.11 ±0.26$\pm$ 0.26± 0.26 94.17±0.01$\pm$ 0.01± 0.01 94.00 ±0.25$\pm$ 0.25± 0.25 93.08 ±0.06$\pm$ 0.06± 0.06
25% 91.46±0.80$\pm$ 0.80± 0.80 89.64±0.76$\pm$ 0.76± 0.76 88.59±1.69$\pm$ 1.69± 1.69 85.14 ±1.07$\pm$ 1.07± 1.07 93.80 ±0.14$\pm$ 0.14± 0.14 93.20±0.19$\pm$ 0.19± 0.19 93.17 ±0.73$\pm$ 0.73± 0.73 92.11 ±0.42$\pm$ 0.42± 0.42
@@
We observe that the accuracies of each submodel for the same percentage of MACs are similar, even for small S architectures with only 25 % of MACs. The accuracies of the full models, highlighted in bold in the top row, are computed using pre-trained models from Zhang et al. (2017) and closely match our experimental results. The performance difference between the BU and TD heuristics is largely insignificant due to the fine-tuning step with extensive training data. Even though the models are overparameterized and in all cases achieve high performance, low-capacity subnetworks S sliced by the bottom-up approach yield minor performance gains. However in the low-data regime, where only few samples per class are available for fine-tuning, the performance differences between BU and TD heuristics are remarkable. Fig. 7 shows the performance when few-shot learning is applied to fine-tune REDS. Our empirical findings confirm a consistently superior performance of the BU heuristic compared to the TD alternative. The differences diminish as more samples are used for fine-tuning (Entezari et al., 2023) and there is barely any difference if full-finetuning is applied using all available training data.
Figure 5: Analysis of the BU knapsack subnetworks from DS-CNN S network, pre-trained on Fashion-MNIST Xiao et al. (2017) and CIFAR10 Krizhevsky et al. (2009). REDS supports vision data domain accuracy without degradation of the pre-trained model.
Table 3: Comparative analysis of REDS and Chowdhery et al. (2019) on MobileNet v1 on VWW pruned down to 0.25%. REDS yields 27% lower peak memory usage for 0.9% lower accuracy.
@@
Figure 6: REDS size S (left) and L (right) BU architectures with ten subnetworks. A larger number of subnetworks does not degrade the accuracy of REDS.
For the MobileNetV1 architecture, we compare REDS to MobileNet v1 (0.25x) using peak memory constraint of 250 KB and the test accuracy previously defined in Chowdhery et al. (2019). REDS knapsack BU heuristic achieves 27% lower peak memory usage for only 0.9% lower accuracy (see Table 3). This validates the efficacy of the knapsack BU approach in constraining peak memory usage within the subnetworks’ architecture. Furthermore, imposing a peak memory usage constraint is essential to facilitate the deployment of the REDS within the confines of the available device RAM (see Fig. 4).
Fig. 3 reports the performance of REDS for accuracy-inference time (left) and accuracy-model size (right) for DS-CNN S and L on the GSC dataset, measured on an Arduino Nano 33 BLE. REDS subnetworks form a strictly superior Pareto frontier: at any given latency or model size, they achieve higher test accuracy than magnitude pruning, SNIP and Slimmable Subnets. These gains stem from REDS’ structured sparsity design, which selects contiguous blocks of filters and encodes each subnetwork via simple slicing points. Thereby eliminating per-inference binary-mask lookups and from its single set of batch-normalization parameters permits compile-time fusion of batch normalization parameters into the convolution weights. In contrast, width pruning methods incur runtime mask-application overhead; and Slimmable Subnets store duplicate batch normalization parameters for each width (i.e., precluding batch normalization fusion), resulting in storage and inference time overheads.
Table 4 presents the Top-1 accuracy of nested ViT-Base subnetworks trained from scratch obtained by considering two MACs constraints (4.60 GMACs and 1.25 GMACs) on ImageNet-1K validation set. Although HydraViT (Haberer et al., 2024) slightly outperforms REDS at the full model configuration (80.45 % vs 79.88 %), REDS achieves superior accuracy at both the medium (78.42 % vs 78.40 % and 77.79 %) and low (68.21 % vs 67.34 % and 66.64 %) subnetworks. This improvement is due to REDS’s iterative ViT knapsack formulation, which allocates attention heads and projection dimensions to maximize the importance score within fixed MACs constraints. In contrast, HydraViT’s fixed head-partitioning and SortedNet’s uniform scaling do not adaptively prioritize the most informative components as MACs constraints become limited. Consequently, REDS achieves consistently higher accuracy–compute trade-offs without altering the underlying contiguous storage of subnetwork parameters.
Fig. 8 visualizes differences between REDS structures obtained with the BU and TD heuristics when using the same initially pre-trained DS-CNN models. Colored bars show the number of filters kept by each nested model in every layer by BU. Overlaid black bars correspond to the number of kept TD filters for the same layer and the number of MACs. The plot reveals consistent differences in each layer as we move from parent to child submodels. The distribution of computational units across subnetworks in each layer depends on the model architecture (see supplementary extended appendix C).
Figure 7: DS-CNN S and L few-shots finetuning on GSC Warden (2018) with MACs percentages as constraints. The subnetworks obtained from the BU Knapsack formulation exhibit faster recovery in accuracy compared to the TD Knapsack ones.
Table 4: Comparison of Top-1 accuracy for nested ViT-Base subnetworks trained from scratch under two MAC constraints. The REDS subnetworks, derived from the ViT-Base TD knapsack formulation, match the MAC counts of the standard N-head ViT-Base models (4.60 and 1.25 GMACs) while employing larger embedding dimensions. For the full model (i.e., 17.56 GMACs), HydraViT achieves 80.45% accuracy, surpassing the 79.88% of REDS. However, at the medium (4.40 GMACs) and low (1.24 GMACs) MACs constraints, REDS outperforms both HydraViT and SortedNet.
Method Dim MACs [G] Acc. [%]
HydraViT (Haberer et al., 2024) 768 17.56 80.45
384 4.60 78.40
192 1.25 67.34
REDS 768 17.56 79.88
512 4.40 78.42
256 1.24 68.21
SortedNet (Valipour et al., 2023) 768 17.56 79.71
384 4.60 77.79
192 1.25 66.64
@@ 5 REDS Optimization for Caches
Embedded ML frameworks like TFLite typically store model weight matrices in row-major order. This means that each row of a weight matrix is stored contiguously in memory. Without the use of subnetworks these weights are also accessed in a contiguous fashion. However, when using subnetworks for model inference, some neurons and their weights are omitted, resulting in non-contiguous memory access. This effect is illustrated in Figure 1. We now show how by adapting the computational graph at compile time, we are able to optimize the computation of REDS subnetworks for devices with a cache memory architecture.
@@ 5.1 Row-major and column-major stores
[…]
@@ 5.2 Optimizing REDS computational graph
The optimization proposed above can be implemented by assuring that each matrix multiplication within a computational graph of each submodel follows the cache optimized access pattern and the corresponding weight matrices are stored in column-major order in memory. We now show the effect of using this optimized computational graph by benchmarking different matrix multiplications. To this end, we use a Raspberry Pi Pico board, which features a RP2040 chip based on a dual-core Arm Cortex-M0+ processor architecture with 264kB RAM and 16MB off-chip flash memory RP2040 (2023). The chip also features execute-in-place (XIP) support, such that the flash memory can be treated like internal memory. These accesses to flash are cached with a 16kB, two way set-associative cache with 8 byte wide cache lines.
Fig. 9 shows the cache effect on the duration of matrix multiplications when we use the optimized computation graph compared to the basic one, i.e., the default way where W is stored in row-major mode and we calculate $X^{T}\cdot W$. We benchmark the matrix multiplication at different splits and use 25%, 50%, 75% and 100% of the neurons in $W$. In Fig. 9 (left) we show the duration of a matrix multiplication with differently shaped weights $W$ times an input matrix $X$ of shape 256×4256$\times$ 4256 × 4, where the weights and inputs are 32 Bit wide floating points. Similarly, Fig. 9 (right) shows the duration of a weights matrix multiplication of size 512×256512$\times$ 256512 × 256 with the same $X$, but now the weights and inputs are unsigned integers with different widths, i.e., 8, 16 and 32 Bit. We observe in all cases a clear speed-up in matrix multiplication when comparing the optimized and the basic computation approach, which is also evident from Table 5. For all test cases the cache-hit rate is above 97% when using optimized matrix multiplications. Additionally, we see a notable difference between floating point and integer based matrices. This difference is due to the fact that the RP2040 does not feature any floating point arithmetic support and, thus, the cache effect is diminished due to longer duration of basic multiplications. Finally, there is no notable difference in speed-up and cache-hit rate between the different splits.
@@
Figure 11: DS-CNN S (left) and L (right) BU comparison of REDS, early exit linear classifiers and magnitude width multiplier pruning subnetworks.
@@ 6 REDS on Mobile and IoT
We evaluate REDS models on two mobile phones, specifically Xiaomi Redmi Note 9 Pro and Google Pixel 6, and on two IoT devices, namely Arduino Nano 33 BLE Sense and Infineon CY8CKIT-062S2. For the mobile phone evaluation, we employed the Google TFLite benchmarking tool, which measures the model’s inference time on Android-based mobile devices. On IoT devices we deployed and evaluated the models using the Edge Impulse (Hymel et al., 2022) platform. We report the number of parameters, accuracy, and latency on the mobiles and on the IoT devices for the full architectures (100% MACs) and the subnetworks architecture (75%, 50% and 25% MACs) found by the BU knapsack solvers. The obtained results are reported in Fig. 10 for the DNN, CNN and DS-CNN architectures S (top row) and L (bottom row) on GSC. All results are averages over three runs. TFLite and TFLMicro lack support for runtime adaptation of model weight tensors. We extended the TFLMicro framework to support REDS out of the box and measure on-device inference and submodel switching times. The code is publicly available [^2] Link: https://github.com/aschesklave/TFlite-Micro-Cutting.
@@
The first left-most column of plots in Fig. 10 shows the relationship between the subnetworks’ MACs and the number of model parameters. All curves are close to linear, yet the parameters of these linear relationships are architecture-specific. The DS-CNN models have the lowest number of parameters, thanks to their more sophisticated architecture, which is more efficient on our dataset compared to the standard convolutional and dense networks. DS-CNN models also yield better accuracy, even for only 25% of MACs, while DNN models perform worst, as can be observed in the second column of plots. This result can be attributed to the successful generalization of the iterative knapsack problem to support depth-wise convolutions Sec. 3.4.2.
Table 6: Energy consumption of REDS size S architectures measured by the Power Profiler Kit (PPK2) on Nordic nRF52840. The results are obtained by running an inference pass through each subnetwork with input data of batch size one, while recording the corresponding inference current.
MACs DNN CNN DS-CNN
[mJ] [mJ] [mJ]
100% 0.18 90.61 61
75% 0.15 86.01 44.03
50% 0.07 50.3 33.81
25% 0.05 44.81 20.44
We compare REDS to several works. We consider model optimization by pruning an equal share of computational blocks in each layer (Tan and Le, 2019) (i.e., width multiplier pruning). We use the block selection strategy based on the magnitude of the weight (i.e., magnitude pruning) used in Cai et al. (2020), and a random selection. Additionally, we compare to Slimmable Subnets (Yu et al., 2019), where each layer stores one extra batch norm statistic for each subnetwork, also obtained by width multiplier pruning. We did not compare REDS to DRESS or NestDNN because neither method’s source code is publicly available, nor do their published descriptions furnish sufficient implementation details to enable reproducible evaluation.
We first consider fixed MACs percentages for presentation simplicity between subnetworks’ methods for DS-CNN S. REDS knapsack solution outperforms all three approaches, achieving higher accuracy. Moreover, we evaluate REDS against early exit linear classifiers (Scardapane et al., 2020) and magnitude-based width multiplier pruning on the DS-CNN S and L architectures concerning MACs and test set accuracy (see Fig. 11). The subnetworks found by the REDS knapsack depth-wise formulation demonstrate superior performance over both methods, particularly within the domain of lower parameter configurations, specifically the S size.
@@
The last two plots on the right show REDS performance on mobile and IoT devices. We observe a difference of three orders of magnitude in the inference times between the two categories. DNN models perform best thanks to the more optimized algorithm and libraries for matrix-matrix multiplication (Goto and Geijn, 2008). All models show a linear relationship between the percentage of MACs and inference time. This empirical evidence validates MACs as a robust predictor of model latency, extending its applicability to mobile devices.
@@
7 Conclusion
This paper presents Resource-Efficient Deep Subnetworks (REDS), a novel approach to adapt deep neural networks to dynamic resource constraints typical for mobile and IoT devices. REDS formulates the subnetworks architectures search as an iterative knapsack problem, taking into account the dependencies between layers, MACs and peak memory usage. Moreover, REDS employs neuron permutation invariance to facilitate adaptation to variable resource availability without compromising the models’ accuracy or adding runtime overhead. Notably, REDS ensures that the subnetworks’ weights are stored in contiguous memory, enhancing cache optimization and compatibility with hardware-specific enhancements like vector instructions. Experimental evaluation on seven benchmark architectures demonstrates the effectiveness of REDS on mobile and IoT devices and superior performance of current pruning and neural architecture search state of the art methods. We perform a theoretical analysis of the knapsack solution space and prove the worst-case performance bounds for the two heuristic algorithms.
There are several directions to extend this work. First, we have not tested the efficiency of REDS used explicitly as neural architecture search to discover neural network architectures for large language models and multimodal models. Second, the support of REDS optimization for caches in convolution models has not been explored and should be addressed in the future to support further advanced architectures. Third, scaling REDS solver to consider more constraints, e.g., energy, is left for future work. We hope this work sparks interest in exploiting neural network symmetries to advance adaptive neural network deployment at the edge.
Acknowledgments
The authors are grateful to Markus Gallacher for his support with the energy efficiency analysis of REDS. Christopher Hinterer and Julian Rudolf contributed to extending the TFL Micro framework to support REDS on Arduino Nano 33 BLE Sense. This research was funded in part by the Austrian Science Fund (FWF) within the DENISE doctoral school (grant number DFH 5). The results presented in this paper were computed using the computational resources of the Vienna Scientific Cluster (VSC) and the HLR resources of the Zentralen Informatikdienstes of Graz University of Technology.